# VITAL: More Understandable Feature Visualization through Distribution Alignment and Relevant Information Flow

### v1

@@ 1 Introduction
Deep neural networks have achieved remarkable success across science and industry. Their reasoning process is, however, inherently opaque. Different lines of research in Explainability of Machine Learning have been addressing these issues as understanding of how a neural network arrives at a decision is crucial, especially in safety-critical domains. To understand which input features are used for a decision, both post-hoc, as well as built-in methodologies have been proposed, which give information about input-output relations in the form of attribution or relevance scores [18, 36, 39, 4]. Yet, these offer only a partial understanding of a model’s decision-making process. In the scope of Mechanistic Interpretability [33, 3], experts are interested in finding explanations that uncover what the model internally learns. In this line of research, methods that discover paths of information flow – so-called circuits – have been proposed [11, 6, 37], along with approaches to discover network components, such as groups of neurons, that together encode a concept [10]. While these methods reveal where information is propagated towards a particular decision, they fall short in identifying which information is encoded. Feature visualization (FV) bridges this gap by generating images that correspond to what triggers a strong response in a given neuron or group of neurons. This is traditionally achieved by optimizing for an image that maximizes the response of the selected neuron(s) through gradient ascent. Complemented with circuits, FV can thus help represent the internal reasoning process into human-understandable terms.
@@ 3 VITAL Framework
Existing FV methods suffer from overly repetitive patterns, artificial, and irrelevant features. To tackle these issues, we suggest a new paradigm focused on optimizing an alignment between features of the learned image and features of real images. By using distributions of activations across different layers, we ensure that both coarse- as well as fine-grained information in our FV corresponds to a typical input (sec. 3.2). This discourages the use of artificial features and repetitive patterns, which typically spike the activation in classic approaches. We further suggest to take into account relevance scores, matching the distributions of activations weighted by their relevance to the target, which discourages features in the FV that are irrelevant for the target neuron (sec. 3.3). This becomes particularly important for the visualization of inner neurons. For example, while grass should be irrelevant for a bird’s beak, it might still co-activate channels across multiple bird images, leading to misleading interpretations. Incorporating LRP eliminates irrelevant correlations in activation maps, ensuring that only the features contributing to the target neuron’s activation are visualized. We provide an example in Fig. 2. We provide further details about our approach in supp. Appendix A.
@@ 3.2 Feature Distribution Matching
where $\text{dist}(\cdot)$ indicates the distribution of the activations.
@@ 3.2.2 Selection of Reference Images
Typically, one would select a set of images from data that is close to the distribution that the model of interest was trained for, e.g. training that the target neuron of interest is highly activated for. For class neurons, we directly use a random selection of images from that class. For intermediate neurons, we select the top-$k$ patches (from distinct images) relevant for this neuron, following Fel et al. [10]. In brief, we crop and resize training images into patches and score each based on the global average pooled activation across spatial dimensions, keeping the top-$k$ patches from unique images. These define our reference $\mathcal{X}_{ref}$ for feature alignment. We provide results on the effect of the sampling in the supp. Sec. C.4.
@@ 3.3 Incorporating Relevance Scores
We thus suggest matching feature distributions based on activation times relevance, i.e., match features $A^{(l^{\prime})}(x)\odot R_{n}^{(l^{\prime})}(x),$ where $R_{n}$ are the relevance scores for the target neuron $n$ under study. These relevance scores give weight to each neuron (or pixel of a channel) describing how relevant the neuron is for target neuron $n$ on the given sample. They are hence of the same dimension as the activations, $\odot$ thus indicating the Hadamard product. In principle, any attribution approach can be used as plug-in replacement to obtain relevance scores $R$. For the remainder of the paper we use LRP, one of the most prominent and successful attribution methods, and provide a comparative result for Guided Backpropagation [38] in supp. Sec. C.5. Besides intermediate neurons, we also applied relevance scores to class neurons in supp. Sec. B.2.
@@ 3.4 Further Considerations
Transparency Map. Irrelevant areas of the generated image stay mostly unchanged during the optimization, essentially representing noise. Analogous to Fel et al. [9], we suggest using transparency maps based on the importance of the image location during optimization to show relevant image parts only. In brief, we accumulate the gradients of our loss (Eq. 2) across each step in the optimization. We thus ensure the identification of the areas that have been most attended to by the network during the generation of the image. In the supp. Sec. C.3, we illustrate the effect of the transparency map for example visualizations.
@@
where $\mathcal{X}_{ref}$ are the set of real reference images, $\mathcal{L}_{\text{SM}}\left(x^{*},\mathcal{X}_{ref}\right)$ is the loss term as specified in Eq. 2 summed over different layers. We ablate on different hyperparameter settings to show their impact on the final visualization in the supp. Sec. C.2.
Choice of Layers for Alignment. While agnostic to the architecture, it is more efficient to focus on aligning features from only few, important layers. To decide on which layers to consider, we conducted a small ablation study (see supp. Sec. C.1) to examine the effects of layers on the visualizations. In a ResNet50, we find that only optimizing for feature alignment of first and last block output is sufficient to produce high-quality images, already capturing the important low- respectively high-level features in the image.
Visualization of Concepts. As part of Mechanistic Interpretability, people are interested in finding concept-based explanations of model behavior. These concepts might be feature directions encoded through multiple neurons in a layer, for example, discovered by CRAFT [10]. In VITAL, we modify the initialization of relevances of target neurons in LRP to reflect the weights given by the feature direction. Through this modification, VITAL can give meaning, to these feature directions. We provide technical details and results of this approach in the supp. Sec. B.3.
@@
Figure 3: Example class visualizations. We show ImageNet class visualizations for (a) ResNet50 and (b) ViT-L-32.MACO and Fourier-based FV (top rows) produce repetitive, hard-to-interpret patterns. DeepInversion (3rd row) improves interpretability but introduces artifacts and lacks ViT compatibility. VITAL arguably yields much more interpretable visualizations. For additional results, including failure cases, we refer to supp. Appendix B.
@@ 4 Experimental Results
To evaluate VITAL, we compare against Fourier [28] as standard baseline, DeepInversion (DeepInv) [44] as state-of-the-art class visualization, and MACO [9], which is generally considered state-of-the-art for both class and inner neuron visualizations. We implemented VITAL in PyTorch and compared on ResNet50 [14], DenseNet121 [15], ConvNext-base [20], ViT-L-16 [8], ViT-L-32 [8], both qualitatively and quantitatively. All architectures were pre-trained on Imagenet. We further conducted human user studies assessing human interpretability of the resulting visualizations. We provide details on implementation and setup of competing methods with more results in the supp. Appendices A and B, respectively.
@@ 4.1 Qualitative Results
Visually, there is a stark difference between class visualizations of VITAL and existing work (see Fig. 3 for ResNet50 and ViT-L-32 and supp. Appendix B for all architectures). Where most existing methods, including MACO and Fourier-based activation maximization, show highly repetitive patterns, our approach provides much cleaner, more understandable representations of the class. While DeepInversion shows less repetition of patterns, there are often irrelevant background features, as in the case of the candle, or visual artifacts, as the blue and yellow color for the dog, dominating the image. Yet, none of the methods is flawless: All approaches, including VITAL, show difficulties with complex spatial arrangements (supp. Sec. B.6). Still, it is usually understandable what class is represented by VITAL. This becomes even more evident when comparing methods across more complex architectures (e.g., ViT, ConvNeXt), where most approaches struggle to remain interpretable. In contrast, VITAL consistently produces representations where the object and its class remain clear, highlighting its robustness and effectiveness.
@@ 4.1.1 Analysis of the Embedding Space
To understand how realistic and distinct the visualizations are, we compare FVs of five different dog classes to original images in terms of their penultimate layer embedding. We show a 2D low-dimensional representation of the embeddings of original and generated images using tSNE [42] in Fig. 5. We find that the real images form five clusters neatly separating the classes. We further observe that fourier-based approaches with their repetitive, often unrecognizable patterns are collapsed on a single point, away from any real images. While DeepInversion does fall into the correct clusters, it is hard to recognize a dog in the generated images. VITAL is the only method that consistently falls into the center of the cluster, providing distinct dog features we can recognize as known dog breeds. We extend our results for other architectures in supp. Sec. B.5.
@@ 4.1.2 Small Circuits
With our approach, it is possible to investigate what inner neurons in a network encode. By extracting the top-3 channel in the last hidden layer most relevant to a specific class ranked by relevance through LRP, we can visualize small mechanistic circuits that show how specific patterns drive a class decision (see Fig. 4). We note that except for DeepInversion, all approaches can be used to visualize inner neurons. For illustrative purposes, we selected MACO as it is the best-performing existing approach. For example, we see that the most relevant, distinguishing features for a zebra classification are the stripes and texture of their fur. For dogs, there are leg, ear, and specifically colored fur detectors evident in the model. Such specific coloring is necessary to distinguish the plenty of dog breeds available as classes in ImageNet. We also observe signs of overfitting, where the pineapple class is associated with neurons that detect groups of fruits, many of the training images show such a picture. While MACO also yields inner neuron visualizations, for the animal classes they are often inherently hard to understand, which we will also see in quantitative terms next.
@@ 4.2 Quantitative Results
For quantitative evaluation, we consider the classification accuracy and the FID score [34] on the respected model, as well as CLIP zero-shot prediction performance [31] in Sec. 4 (see supp. Sec. A.3 for the extended version). We also provide quantitative results for the intermediate neurons with the metrics proposed by [17] in supp. Sec. B.1.
Accuracy. Visualizing each class neuron and then predicting the class of the visualization, we measure how accurate of a description of the target these visualizations are (see Tab. 4). While existing work optimizes for class logits, perhaps surprisingly, the achieved accuracy for all methods except DeepInversion is low. Only VITAL and DeepInversion are able to produce accurate descriptions, achieving virtually 100% accuracy. Investigating further, we observe that other methods including MACO seem to produce irrelevant features that mislead the model, producing predictions unrelated to the original class. We give two examples of prediction in supp. Sec. B.4.
FID Score. For all architectures, to quantify the similarity between the feature distributions of visualizations to that of natural images within the same object category in Sec. 4, we use FID scores [34] computed on features extracted at the penultimate layer of ResNet50. While this metric is aligned with our optimization, it is surprising to see the extent of the difference to most other methods, showing over a magnitude better score. Similar to VITAL, DeepInversion also considers feature statistics, namely batch normalization statistics. While in the optimal case reaching the best performance, the choice of batch size can have a high impact on visualization quality and the FID score.
CLIP Zero-shot Prediction. So far, the considered metrics were based on the model for which visualizations were generated. To understand how a different model perceives the visual features presented in a FV, we consider a pretrained CLIP ViT-B/32 model and query it with templates to generate class labels for the FVs (see Appendix for more details). Based on these answers, we compute a zero-shot prediction score for all classes and each method. We provide scores including those of original (correctly classified) ImageNet images in Tab. 4. As before, we see that DeepInversion outperforms existing work, with standard baselines having overall poor performance even when considering Top-5 accuracy. Furthermore, evaluated on this separate model, VITAL surpasses the performance of all existing work by a wide margin, coming close to that of real data images.
@@ 4.3 Human Interpretability Study
We provide an overview of the results of our study in Fig. 6. We observe that our method performs favorable across all three tasks compared to other methods, showing an order of magnitude more high-scoring visualizations (scores 4-5) in both (1) and (2), and showing much higher similarity to the correct class on task (3). In particular, for this challenging task that provides a zero-shot estimate of the interpretability of FVs, the 5% quantile, i.e., the ”worse”-performing visualizations, are still close to a similarity of .5.5.5.5 of VITAL. Other methods show only $0.25$ similarity for the same quantile. The same stark difference of more than $0.2$ in similarity score is evident in the median, showing that VITAL provides more human-understandable visualizations of what neurons encode. We proved more fine-grained analysis, population statistics, and details in supp. Appendix D. Additionally, we conducted a human user study following the protocol of [48, 9], where VITAL performs best (supp. Appendix D). The study design provides a simple option-based evaluation similar to our setting (2).
@@ 5 Discussion
Our experiments show that VITAL yields better understandable visualizations, more accurately capturing the distinct information encoded by a neuron, outperforming its most recent competitor MACO [9] in terms of classification accuracy for the generated visualizations on the target model. While DeepInversion shows similar performance on this metric, this comes to little surprise, as their approach explicitly maximizes the logit class score of that (same) model. When asking a different neural network, such as a CLIP model, to label the images, however, our approach outperforms all competitors by a wide margin. Similarly, two human user studies we conducted provide evidence that visualizations of VITAL are more understandable than those of competitors in describing given labels or images.
@@
Realism. In principle, our visualizations, as well as those of other methods, do not reach photo-realism. VITAL visualizations sometimes resemble the style of paintings of Édouard Manet, which, however, does not limit their interpretability: Similar to a painting, it is understandable what is expressed, which is also supported by the user studies. Yet, it would be interesting to study how to reach photo-realism in the future. While generative approaches such as diffusion models is an interesting avenue, it requires further research on how to avoid generative-model specific features in the generated image, as these erode the faithfulness of the visualizations.
@@
Evaluation. For our evaluation, we considered a diverse set of metrics to evaluate our framework against the state-of-the-art as comprehensively as possible, leveraging what is common in the literature and adding further metrics, such as through CLIP prompting. Unfortunately, in feature visualization, but also in Explainability research more broadly, exhaustive benchmarks and clearly defined metrics are actively discussed, but currently still amiss. Through the additional human user studies, we complement the metrics with an evaluation of our original goal of more understandable feature visualizations. As with all human user studies, its design can always be debated. More refined study settings and including larger cohorts would perhaps make for an interesting study on its own.
@@
Distribution Matching. Standard approaches to distribution matching include matching statistics, such as mean and variances, or considering prior distributions [16, 21], but these oversimplify the complexity of real images. We observed a diverse set of activation distributions that are hard to capture with a (simple) distribution prior or first or second moment statistics [47]. Hence, directly optimizing an alignment between the empirical distributions showed the most promise, with sort matching [32] providing an efficient approach that we can scale easily to multiple layers of modern networks. In principle, our approach is agnostic to the type of distribution matching algorithm.
Future work. Finally, there is always room for improvement. Especially both human user studies showed that there is still more to gain, the generated images are not consistently scored as perfectly reflecting the given class or image. Our method provides a substantial improvement over existing work and we anticipate it of great use in downstream applications, such as understanding networks in the medical domain, studying what neurons encode after knowledge transfer, or how pruning affects the representations.

### v2

@@ 1 Introduction
Deep neural networks have achieved remarkable success across science and industry. Their reasoning process is, however, inherently opaque. Different lines of research in Explainability of Machine Learning have been addressing these issues as understanding of how a neural network arrives at a decision is crucial, especially in safety-critical domains. To understand which input features are used for a decision, both post-hoc, as well as built-in methodologies have been proposed, which give information about input-output relations in the form of attribution or relevance scores [20, 38, 41, 6]. Yet, these offer only a partial understanding of a model’s decision-making process. In the scope of Mechanistic Interpretability [35, 5], experts are interested in finding explanations that uncover what the model internally learns. In this line of research, methods that discover paths of information flow – so-called circuits – have been proposed [13, 8, 39], along with approaches to discover network components, such as groups of neurons, that together encode a concept [12]. While these methods reveal where information is propagated towards a particular decision, they fall short in identifying which information is encoded. Feature visualization (FV) bridges this gap by generating images that correspond to what triggers a strong response in a given neuron or group of neurons. This is traditionally achieved by optimizing for an image that maximizes the response of the selected neuron(s) through gradient ascent. Complemented with circuits, FV thus help represent the internal reasoning process in human-understandable terms.
@@ 3 VITAL Framework
Existing FV methods suffer from overly repetitive patterns, artificial, and irrelevant features. To address this, we propose aligning feature distributions between learned and real images across layers, capturing both coarse and fine-grained information (Sec. 3.2). This discourages the use of artificial features and repetitive patterns, which typically spike the activation in classic approaches. We further suggest to take into account relevance scores, matching the distributions of activations weighted by their relevance to the target, which discourages features in the FV that are irrelevant for the target neuron (sec. 3.3). This becomes particularly important for the visualization of inner neurons. For example, while grass should be irrelevant for a bird’s beak, it might still co-activate channels across multiple bird images, which misleads interpretations. Incorporating LRP eliminates irrelevant correlations in activation maps, ensuring that only the features contributing to the target neuron’s activation are visualized. We provide an example in Fig. 2 and further details about our approach in Appendix Appendix A.
@@ 3.2 Feature Distribution Matching
where $\text{dist}(\cdot)$ indicates the distribution of the activations. In Fig. 3, we illustrate the alignment of feature distributions for sample channels and blocks.
@@ 3.2.2 Selection of Reference Images
Typically, the reference set should reflect the distribution relevant to the target neuron. For class neurons, we directly use a random selection of images from that class. For intermediate neurons, we select the top-$k$ patches (from distinct images) relevant for this neuron, following Fel et al. [12]. In brief, we crop and resize training images into patches and score each based on the global average pooled activation across spatial dimensions, keeping the top-$k$ patches from unique images. These define our reference $\mathcal{X}_{ref}$ for feature alignment. We provide additional sampling experiments, including size sensitivity, cluster-based sampling, dataset generalization, and corruption analysis, in the Appendix Sec. C.4.
@@ 3.3 Incorporating Relevance Scores
We thus suggest matching feature distributions based on activation times relevance, i.e., match features $A^{(l^{\prime})}(x)\odot R_{n}^{(l^{\prime})}(x),$ where $R_{n}$ are the relevance scores for the target neuron $n$ under study. These relevance scores give weight to each neuron (or pixel of a channel) describing how relevant the neuron is for target neuron $n$ on the given sample. They are hence of the same dimension as the activations, $\odot$ thus indicating the Hadamard product. In principle, any attribution approach can be used as plug-in replacement to obtain relevance scores $R$. For the remainder of the paper, we use LRP, one of the most prominent attribution methods, and provide a comparative result for Guided Backpropagation [40] in Appendix Sec. C.5. Besides intermediate neurons, we also applied relevance scores to class neurons in Appendix Sec. B.2.
Figure 3: Overview for feature distribution matching. Activations or activation $\times$ relevance scores (AA) are extracted from channels across network blocks for real and optimized images. For feature visualization, the distribution of the generated image $x^{*}$ is optimized to match those of the real reference images $x^{\prime}\in\mathcal{X}_{ref}$.
@@ 3.4 Further Considerations
Transparency Map. Irrelevant areas of the generated image stay mostly unchanged during the optimization, essentially representing noise. Analogous to Fel et al. [11], we suggest using transparency maps based on the importance of the image location during optimization to show relevant image parts only. In brief, we accumulate the gradients of our loss (Eq. 2) across each step in the optimization. We thus ensure the identification of the areas that have been most attended to by the network during the generation of the image. In the Appendix Sec. C.3, we illustrate the effect of the transparency map for example visualizations.
@@
where $\mathcal{X}_{ref}$ are the set of real reference images, $\mathcal{L}_{\text{SM}}\left(x^{*},\mathcal{X}_{ref}\right)$ is the loss term as specified in Eq. 2 summed over different layers. We ablate on different hyperparameter settings to show their impact on the final visualization in the Appendix Sec. C.2.
Choice of Layers for Alignment. While agnostic to the architecture, it is more efficient to focus on aligning features from only few, important layers. To decide on which layers to consider, we conducted a small ablation study (see Appendix Sec. C.1) to examine the effects of layers on the visualizations. In a ResNet50, we find that only optimizing for feature alignment of first and last block output is sufficient to produce high-quality images, already capturing the important low- respectively high-level features in the image.
Visualization of Concepts. As part of Mechanistic Interpretability, people are interested in finding concept-based explanations of model behavior. These concepts might be feature directions encoded through multiple neurons in a layer, for example, discovered by CRAFT [12]. In VITAL, we modify the initialization of relevances of target neurons in LRP to reflect the weights given by the feature direction. Through this modification, VITAL can give meaning, to these feature directions. We provide technical details and results of this approach in the Appendix Sec. B.3.
@@
Figure 4: Example class visualizations. We show ImageNet class visualizations for (a) ResNet50 and (b) ViT-L-32.MACO and Fourier-based FV (top rows) produce repetitive, hard-to-interpret patterns. DeepInversion (3rd row) improves interpretability but introduces artifacts and lacks ViT compatibility. VITAL arguably yields much more interpretable visualizations. For additional results, including failure cases, we refer to Appendix Appendix B.
@@ 4 Experimental Results
To evaluate VITAL, we compare against Fourier [30] as standard baseline, DeepInversion (DeepInv) [46] as state-of-the-art class visualization, and MACO [11], which is generally considered state-of-the-art for both class and inner neuron visualizations. We implemented VITAL in PyTorch and compared on ResNet50 [16], DenseNet121 [17], ConvNext-base [22], ViT-L-16 [10], ViT-L-32 [10], both qualitatively and quantitatively. All architectures were pre-trained on Imagenet. We further conducted human user studies assessing human interpretability of the resulting visualizations. We provide details on implementation and setup of competing methods with more results in the Appendix Appendices A and B, respectively.
@@ 4.1 Qualitative Results
Visually, there is a stark difference between class visualizations of VITAL and existing work (see Fig. 4 for ResNet50 and ViT-L-32 and Appendix Appendix B for all architectures). Where most existing methods, including MACO and Fourier-based activation maximization, show highly repetitive patterns, our approach provides much cleaner, more understandable representations of the class. While DeepInversion shows less repetition of patterns, there are often irrelevant background features, as in the case of the candle, or visual artifacts, as the blue and yellow color for the dog, dominating the image. Yet, we acknowledge that none of the methods is flawless: All approaches, including VITAL, show difficulties with complex spatial arrangements (Appendix Sec. B.6). Still, it is usually understandable what class is represented by VITAL. This becomes even more evident when comparing methods across more complex architectures (e.g., ViT, ConvNeXt), where most approaches struggle to remain interpretable. In contrast, VITAL consistently produces representations where the object and its class remain clear, highlighting its robustness and effectiveness.
@@ 4.1.1 Analysis of the Embedding Space
To understand how realistic and distinct the visualizations are, we compare FVs of five different dog classes to original images in terms of their penultimate layer embedding. We show a 2D low-dimensional representation of the embeddings of original and generated images using tSNE [44] in Fig. 6. We find that the real images form five clusters neatly separating the classes. We further observe that fourier-based approaches with their repetitive, often unrecognizable patterns are collapsed on a single point, away from any real images. While DeepInversion does fall into the correct clusters, it is hard to recognize a dog in the generated images. VITAL is the only method that consistently falls into the center of the cluster, providing distinct dog features we can recognize as known dog breeds. We extend our results for other architectures in Appendix Sec. B.5.
@@ 4.1.2 Small Circuits
Our method enables interpreting what inner neurons encode. By selecting the top-3 channels in the last hidden layer most relevant to a class (via LRP), we can visualize local circuits that reveal class-specific patterns (Fig. 5). While most methods, except DeepInversion, support inner neuron visualization, we compare with MACO as the strongest baseline. For example, we see that the most relevant, distinguishing features for a zebra classification are the stripes and texture of their fur. For dogs, there are leg, ear, and specifically colored fur detectors evident in the model. Such specific coloring is necessary to distinguish the plenty of dog breeds available as classes in ImageNet. We also observe signs of overfitting, where the pineapple class is associated with neurons that detect groups of fruits, many of the training images show such a picture. MACO yields less interpretable results, particularly for animals, as confirmed in the next section.
@@ 4.2 Quantitative Results
For quantitative evaluation, we consider the classification accuracy and the FID score [36] on the respected model, as well as CLIP zero-shot prediction performance [33] in Tab. 1 (see Appendix Tab. 2 for the extended version). We also provide quantitative results for the intermediate neurons with the metrics proposed by [19] in Appendix Sec. B.1.
Accuracy. To assess how well visualizations describe their target, we classify each class neuron’s visualization (see Tab. 1). Despite explicitly optimizing class logits, most existing methods perform poorly. Only VITAL and DeepInversion are able to produce accurate descriptions, achieving virtually 100% accuracy. Investigating further, we observe that other methods including MACO seem to produce irrelevant features that mislead the model, producing predictions unrelated to the original class. We give two examples of prediction in Appendix Sec. B.4.
FID Score. For all architectures, to quantify visualization realism, we compute FID scores [36] using features from ResNet50’s penultimate layer (Tab. 1). While aligned with our objective, the performance gap to other methods is substantial, with VITAL achieving over an order of magnitude better scores. DeepInversion also leverages feature statistics but is highly sensitive to batch size, which impacts both FID and visual quality.
CLIP Zero-shot Prediction. So far, previous metrics relied on the model used for visualization. To assess general perceptual quality, we use a pretrained CLIP ViT-B/32 model to classify FVs via template-based prompts (see Appendix Sec. A.2 for more details). We report zero-shot scores in Tab. 1, including original (correctly classified) ImageNet images for reference. As before, we see that DeepInversion outperforms existing work, with standard baselines having overall poor performance even when considering Top-5 accuracy. Notably, VITAL outperforms all others by a wide margin, approaching real image accuracy.
Figure 6: t-SNE projection of embedding. We show a low-dimensional tSNE embedding of the features at the penultimate layer for five dog breeds indicated by color. Transparent circles are original training images and FVs are indicated by symbols:
@@ 4.3 Human Interpretability Study
We provide an overview of the results of our study in Fig. 7. We observe that our method performs favorable across all three tasks compared to other methods, showing an order of magnitude more high-scoring visualizations (scores 4-5) in both (1) and (2), and showing much higher similarity to the correct class on task (3). In particular, for this challenging task that provides a zero-shot estimate of the interpretability of FVs, the 5% quantile, i.e., the ”worse”-performing visualizations, are still close to a similarity of .5.5 of VITAL. Other methods show only $0.25$ similarity for the same quantile. The same stark difference of more than $0.2$ in median similarity score shows that VITAL provides more human-understandable visualizations of what neurons encode. We provide more fine-grained analysis, population statistics, and details in Appendix Appendix D. We further conducted a human user study following the protocol of [50, 11], which is a simple option-based evaluation similar to setting (2), where VITAL performs best (Appendix Appendix D).
@@ 5 Discussion
Our experiments show that VITAL produces more understandable visualizations that better capture the unique information encoded by each neuron, outperforming MACO [11], the most recent method in classification accuracy on the target model. While DeepInversion performs similarly, this is expected since it directly optimizes the model’s logit score. When evaluated by a different model such as CLIP, VITAL surpasses all baselines by a wide margin. Two human user studies further confirm that VITAL yields more interpretable visualizations in describing given labels or images.
@@
Realism. In principle, our visualizations, as well as those of other methods, do not reach photo-realism. VITAL visualizations sometimes resemble the style of paintings of Édouard Manet, which, however, does not limit their interpretability: Similar to a painting, it is understandable what is expressed, which is also supported by the user studies. Yet, it would be interesting to study how to reach photo-realism in the future. Generative models such as diffusion offer a promising direction, though further research is needed to prevent the introduction of model-specific artifacts that reduce faithfulness to the target model.
@@
Evaluation. We evaluate our framework using a broad set of metrics, combining standard benchmarks with additional measures like CLIP prompting. Unfortunately, in feature visualization, more broadly in Explainability research, exhaustive benchmarks and clearly defined metrics are actively discussed, but currently still amiss. Through the additional human user studies, we complement the metrics with an evaluation of our original goal of more understandable feature visualizations. As with any user study, the design can be debated, and more refined setups with larger cohorts offer a promising direction for future work.
@@
Distribution Matching. Standard approaches to distribution matching include matching statistics, such as mean and variances, or considering prior distributions [18, 23], but these oversimplify the complexity of real images. Activation distributions are often too diverse for such approximations [49]. Hence, directly optimizing an alignment between the empirical distributions showed the most promise, with sort matching [34] as an efficient approach that easily scales to modern networks, but our approach is agnostic to the type of distribution matching algorithm.
Future work. VITAL provides a substantial improvement over existing work and we anticipate it of great use in downstream applications, such as understanding networks in the medical domain, studying what neurons encode after knowledge transfer, or how pruning affects the representations. Future directions include automatic selection of neurons or neuron sets for circuit visualization, for example using CRP [3], integration of ViT-specific visualization via AttnLRP [4], and extending our analysis to CLIP-based ViT models to better understand multimodal representations.